OpenStudy (cenaida):

Use the Pythagorean Theorem to find the length of the missing side of the right triangle with both legs equal to √2. a right triangle with legs as square root of 2 and square root of 2. A. 2 B. 2 √2 C. √5 D. None of the above

OpenStudy (nirmalnema):

its A ...=2

OpenStudy (nirmalnema):

because...\[\sqrt{2}^{2}+\sqrt{2}^{2}= (hyp)^{2}\]

OpenStudy (nirmalnema):

\[\sqrt{4}=2\]

OpenStudy (nirmalnema):

so a is correct

OpenStudy (cenaida):

thanks! :D

OpenStudy (nirmalnema):

^-^